OpenStudy (anonymous):

Scores on a university exam are normally distributed with a mean of 68 and a standard deviation of 9. Using the 68-95-99.7 rule, what percentage of students score below 50?

Directrix (directrix):

z = (50 - 68) / 9 = -18/9 = -2 The score of 50 is located two standard deviation below the mean.

Directrix (directrix):

Looking at the standard normal curve (see attachment), what percentage of area under the curve falls below z = - 2. @B_Sosa21

Directrix (directrix):

Change to a percent.